OpenStudy (anonymous):

WILL FAN AND MEDAL! :) All I need to know is how to set up the equation- Do I change the percentage to 0.035, 0.35 or leave it as 3.5? And do I add 1? Aidan has $7565 in his checking account. He invests it in an account that earns 3.5% interest compounded continuously. What is the total amount of his investment after 3 years?

OpenStudy (yb1996):

Each time you use a percent in an equation, you should convert it into a decimal. So 0.035 will be correct.

OpenStudy (anonymous):

so it will turn into A=7565(1+0.035)^3?

OpenStudy (yb1996):

The equation you will use will be P=e^(rt)

OpenStudy (yb1996):

Sorry, A = Pe^(rt)

OpenStudy (anonymous):

how do you set that up?

OpenStudy (yb1996):

This website has a nice explanation on how to use that equation. Read it and if you have any questions then let me know: http://www.mathwarehouse.com/compound-interest/continuously-compounded-interest.php

OpenStudy (anonymous):

so A=7565e^(0.035)(3)

OpenStudy (yb1996):

Yes!

OpenStudy (anonymous):

what does 3 represent?

OpenStudy (yb1996):

3 years

OpenStudy (anonymous):

i meant e, sorry

OpenStudy (anonymous):

so right now i have 7565e^0.105

OpenStudy (yb1996):

e is a constant. It's approximately 2.7. But you should be able to find it on your calculator.

OpenStudy (anonymous):

how do i find that using a calculator

OpenStudy (yb1996):

do you have a scientific calculator with you?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wait, I got it

OpenStudy (yb1996):

Alright, I was going to ask you which calculator you had since all calculators have different ways of inputting e, but that makes my life a lot easier...lol

OpenStudy (yb1996):

you should close the question so that it goes off the list